OpenStudy (anonymous):

Use rules of logarithms to expand log (x+3)^2 / (x-2)(x^2+5)^4

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

That entire thing is under a log, yes?

OpenStudy (anonymous):

Yes, the log symbol is before the division

terenzreignz (terenzreignz):

Okay, well, here's the gist of it... logs turn products into sums and quotients into differences... Here's how it works: \[\Large \log\frac{ab}{cd}=\log(a)+\log(b)-\log(c)-\log(d)\] got it so far? every factor in the numerator gets a plus-sign, every factor in the denominator gets a minus sign. Use that to your advantage, and tell me what you get...

OpenStudy (anonymous):

hmmm but how do I get ab/cd with the exponents

terenzreignz (terenzreignz):

Don't worry about those for now, treat those, say, (x+3)^2 as a single entity/factor for now, we'll deal with those exponents AFTERWARDS :>

OpenStudy (anonymous):

So log(x+3) + log(x+3) - log(x-2) - log(x^2+5)?

terenzreignz (terenzreignz):

You forgot the ^4 on the (x^2 + 5) And I just told you to let those exponents alone :P

terenzreignz (terenzreignz):

But I think you got the idea, sort of... you were close... \[\Large \log(x+3)^2 - \log(x-2)-\log(x^2+5)^4\] right?

terenzreignz (terenzreignz):

This is not yet finished, btw... but, do you catch me so far? ^_^

OpenStudy (anonymous):

Yes

terenzreignz (terenzreignz):

Okay, NOW we deal with the exponents... using this simple property of logs: \[\Large \log (t)^\color{red}p= \color{red}p\log(t)\] Got it? Recap: It turns products into sums: It turns quotients into differences: It turns exponents into factors ^_^ As you can see, the power can be 'brought down' to be multiplied to the log instead.

OpenStudy (anonymous):

2log(x+3) - log(x-2) - 4log(x^2+5)?

terenzreignz (terenzreignz):

Perfect :) That's all we can do. Good job ^_^

OpenStudy (anonymous):

So that's my answer expanded!?

terenzreignz (terenzreignz):

Yup.

OpenStudy (anonymous):

It has the 2 in front? Sometimes I know the 2 would go as a little number on the right of the log

terenzreignz (terenzreignz):

Only the log(x+3) would have a 2, since it had an exponent of 2 earlier.

terenzreignz (terenzreignz):

What you did earlier, which was log(x+3) + log(x+3) actually accomplishes the same thing, but please don't make a habit of that, and use th exponent-to-factor shortcut ^_^

OpenStudy (anonymous):

Ok as long as your sure I got the answer, then thanks! Could you help on one more?

terenzreignz (terenzreignz):

We'll see.

OpenStudy (anonymous):

Rewrite using radicals: (a^3/4)^5

terenzreignz (terenzreignz):

an entirely different problem, i see? Well then, use the laws of exponents first... \[\Large (a^m)^n = a^{mn}\]

OpenStudy (anonymous):

a^3/4 * 5

terenzreignz (terenzreignz):

Which is...?

terenzreignz (terenzreignz):

Multiply fractions... elementary... no? \[\Large 5\times \frac34 = ?\]

OpenStudy (anonymous):

3.75

terenzreignz (terenzreignz):

Please keep it as a fraction...

OpenStudy (anonymous):

15/4

terenzreignz (terenzreignz):

Okay, good... \[\LARGE = a^{\frac{15}4}\] now, use this fact: \[\LARGE a^{\frac{\color{red} m}{\color{green} n}}= \sqrt[\color{green} n]{a^\color{red}m} \]

OpenStudy (anonymous):

4sqrt a^15

terenzreignz (terenzreignz):

Okay great... that should do it... though it CAN be simplified further.

OpenStudy (anonymous):

So that isnt my answer?

terenzreignz (terenzreignz):

It does get the job done, though... but ideally, the exponent inside the radical shouldn't be bigger than the index of the radical [4]. So rewrite it as \[\Large \sqrt[4]{a^{15}}=\sqrt[4]{(a^3)^4a^3}\] And I'm sure you can see where this is going...

OpenStudy (anonymous):

So should I leave the 4sqrt a^15 as my answer though?

terenzreignz (terenzreignz):

I don't know... it doesn't really say simplify, does it...

OpenStudy (anonymous):

No it doesn't

terenzreignz (terenzreignz):

Well then, go for it, and see how it plays out ^_^

OpenStudy (anonymous):

haha ok thanks! That is all for now